OpenStudy (anonymous):

Find the values of thetha between 0 and 2 pi that satisfy sin thetha = - square root 3 / 2

OpenStudy (jdoe0001):

got your Unit Circle handy?

OpenStudy (anonymous):

sin is negative in third and fourth quadrant.

OpenStudy (jdoe0001):

\(\bf sin(\theta)=-\sqrt{\cfrac{3}{2}}\qquad \qquad 0<\theta<2\pi \\ \quad \\ sin^{-1}[sin(\theta)]=sin^{-1}\left( -\sqrt{\cfrac{3}{2}} \right)\implies \theta=sin^{-1}\left( -\sqrt{\cfrac{3}{2}}\right)\)

OpenStudy (anonymous):

yes i do

OpenStudy (jdoe0001):

so check your unit circle for "those angles" within that constrained range

OpenStudy (jdoe0001):

and keep in mind what surjithayer pointed out you want a negative sine, and that means, the 3rd and 4th quadrants

OpenStudy (anonymous):

Thank you so much! Both of you! :)

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

\[\sin \theta=-\frac{ \sqrt{3} }{ 2 }=-\sin \frac{ \pi }{ 3 }=\sin \left( \pi+\frac{ \pi }{ 3 } \right),\sin \left( 2\pi-\frac{ \pi }{ 3} \right)\]

OpenStudy (anonymous):

if it is\[\sin \theta=-\sqrt{\frac{ 3 }{ 2 }}\] then the calculations of jdoe0001 are correct.

OpenStudy (anonymous):

it is \[\sin \theta = -\frac{ \sqrt{3}}{ 2 }\] @surjithayer

OpenStudy (anonymous):

\[\sin \theta=\sin \frac{ 4 \pi }{ 3 },\sin \left( \frac{ 5 \pi }{ 3 } \right),\theta=?\]

OpenStudy (anonymous):

Is it .091?

OpenStudy (anonymous):

no \[\theta=\frac{ 4 \pi }{ 3 },\frac{ 5 \pi }{ 3 }\]

OpenStudy (anonymous):

you have to find the angles